OpenStudy (anonymous):

-15=3/7x SOLVE for the equation

OpenStudy (anonymous):

(-15*7)/3=x

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

is it -15=(3/7)x or -15=3/(7x)

OpenStudy (anonymous):

if you can write it using the equation thing then i can understand it better